# PEPT: Expert Finding Meets Personalized Pre-training

### v1

@@ 5.4. Evaluation Metrics
For evaluating the model’s performance, we use the following three widely employed evaluation metrics:
@@
Overall, the above diverse metrics are used to assess the ranking accuracy of the expert recommended by an algorithm. A higher value for each of these metrics indicates better performance of the model in accurately predicting the most relevant expert.
@@ 6. EXPERIMENTAL RESULTS
Within this section, we showcase the outcomes of our comprehensive experiments and conduct a detailed analysis of PEPT. Specifically, (1) Performance Comparison compares the PEPT with 10 competitive baselines; (2) Ablation Study explores how different parts of the model affect model performance; (3) Pre-trained Expert Representation Analysis analyzes what the pre-trained expert representations benefit the expert finding. (4) Hyper-parameter Analysis analyzes the model performance with different model hyper-parameters. (5) Case Study understand PEPT’s advantages in finding experts more intuitively.
@@ 6.1. Performance Comparison
Table 4 displays the overall results of our proposed model in comparison to other baselines. The experimental findings reveal several following observations:
@@
    Previous methods, such as Doc2Vec and CNTN, exhibit poor performance on almost all datasets. This attributes to the fact that these methods rely on the max or mean operator on experts’ histories to obtain expert features, which fails to consider the varying importance of different historical questions in modeling the interest of an expert.
@@
    In contrast, recent attention-based methods, such as RMRN and PMEF, have achieved better results. The reason is that these methods leverage attention mechanisms to selectively weigh the importance of different historical questions in modeling expert representations. It is worth noting that the PMEF outperforms most of the baselines. This is because PMEF utilizes a multi-view attentive paradigm, enabling it to learn multi-view representations of experts and questions.
@@
    As we can see, our proposed method PEPT surpasses two very recent SOTA work ExpertBert (Liu et al., 2022b) and ExpertPLM (Peng and Liu, 2022). The reason is that the ExpertBert could not capture expert abilities to answer different questions and the ExpertPLM could not capture more fine-grained question-level interest and expertise. And both of them are incapable of modeling personalized expert representations, which can reduce their ability to accurately model different experts.
In summary, the results validate the motivation proposed in Section 1 and reflect that PEPT with the characteristics in the Table 1 can achieve the best performance.
@@ 6.2. Ablation Study
Table 5. Characteristics of Different Variants.
Characteristics Variants Only Em w/o Expertise w/o Interest w/o Personalized PEPT
Expert-level Information $\surd$ $\surd$ $\surd$ $\surd$ $\surd$
Vote Score $\times$ $\times$ $\surd$ $\surd$ $\surd$
ID Embedding $\times$ $\surd$ $\surd$ $\times$ $\surd$
MLM $\surd$ $\surd$ $\surd$ $\surd$ $\surd$
Vote-oriented Task $\times$ $\times$ $\surd$ $\surd$ $\surd$
Question-level MLM $\times$ $\surd$ $\times$ $\surd$ $\surd$
To further explore how different modules of the pre-training model impact the model performance, we design diverse variants for experimenting (CodeReview and Gis datasets as examples), including:
@@
    w/o Expertise, we incorporate expert interest modeling and personalized information during the pre-training phase, and exclude the vote score information and vote-oriented task from our model;
@@
A general comparison of the four variants is provided in Table 5, and the experimental results are illustrated in Fig 3. Based on these results, we have made the following observations:
Figure 3. Ablation Study.
@@
    It is not surprising that Only Em obtains the worst performance. This is because the Only Em only uses expert-level word sequences (i.e., historical answered questions) and cannot effectively capture the question-level interest and ability required for finding suitable experts;
@@
    w/o Expertise and w/o Interest show varying degrees of decreased model performance. These results strongly reflect the importance of capturing expert interest and capability for finding suitable experts. Interestingly, the performance of w/o Interest is worse than that of w/o Expertise. We argue that interest may be more important than ability. In other words, experts in real life should first consider their interests rather than their abilities before answering questions;
@@
    Personalized information is crucial for modeling experts. As expected, removing the personalized information during pre-training leads to a decrease in model performance, which can be seen in w/o Personalized;
@@
    Our complete model PEPT achieves the highest performance among all the variants. The reason is that the PEPT is equipped with six characteristics in Table 5, which enable personalized pre-training of expert interest and expertise for one specific question.
Overall, the ablation studies provide evidence for our motivation and demonstrate the effectiveness of each component in our proposed model. Combining all key components in the PEPT framework leads to significant improvements in performance.
@@ 6.3.1. Effect of Alleviating Data Sparsity
Figure 4. Impacts of Fine-tuning Data Ratio. GAP = PEPT - PMEF.
Fig 4 shows the results on CodeReview and Electronics datasets. The performance of all models substantially drops when a smaller amount of training data is used. However, our pre-trained model, PEPT, consistently outperforms the best non pre-trained model, PMEF, in all cases. Furthermore, we notice a notable trend in the experimental results: as the amount of training data decreases, the performance gap between our proposed model, PEPT, and the baseline method, PMEF, becomes more pronounced. This observation suggests that the advantage of our pre-trained representations becomes even more pronounced when the training data is limited. This is because our pre-training paradigm enables the model to personalized leverage expert behaviors and vote scores for modeling experts’ interests and expertise. Besides, it could reduce the model’s dependence on training data.
Furthermore, we can find that although ExpertPLM also has the ability to alleviate data sparsity issues, it is still inferior to PEPT. This is because ExpertPLM can only pre-train coarse-grained expert representations, and it still needs sufficient supervised data for fine-tuning to capture specific representations of different experts. Different from that, PEPT can perform personalized expert representation learning during the pre-training stage and capture more fine-grained characteristics for expert finding. Overall, our experimental results suggest that PEPT can mitigate the impact of the data sparsity problem in expert finding.
Figure 5. Comparison with the Original BERT.
Figure 6. Impacts of Different Mask Ratio during Pre-training.
6.3.2. Comparison with the Original BERT
In our proposed method, the personalized expert pre-training paradigm is based on the original pre-trained BERT (i.e., bert-base-uncased). In this part, we further explore the effectiveness of this weight (derived by PEPT) by comparing it with the original BERT in the fine-tuning stage. Specifically, we replace our pre-trained weight with the BERT weight during the fine-tuning stage and compare the resulting performance.
The results are illustrated in Fig 5 (Gis and Biology as examples). Our experimental results indicate that the pre-trained weight obtained through our proposed pre-training paradigm is beneficial to the expert finding. As we noted previously, learning expert representations accurately is crucial for expert finding, as it involves capturing the interests and expertise of experts in answering various questions. However, the bert-base-uncased weight cannot capture these expert characteristics and can only find suitable experts based on text similarity matching. This observation highlights effectiveness of our pre-training paradigm in expert finding.
@@ 6.4. Further Analysis
In this section, we conduct experiments to evaluate the effect of key hyperparameters (including Masked Word Ratio, Masked Question Ratio, Learning Rate inf Fine-tuning) in our approach.
@@ 6.4.1. Masked Word Ratio
The masked word ratio is a critical hyperparameter in our proposed PEPT model during the expert representation pre-training. To investigate the sensitivity of parameters in the pre-training phase, we varied the ratio from $0.05$ to $0.25$ in increments of $0.05$.
The results are presented in Fig 6. All metric results initially improve as the mask ratio increases, reaching the optimal model performance at a certain point, and then gradually declining. If the word mask ratio is set too small, the pre-training process may not capture enough CQA language knowledge and result in reduced performance in the downstream task. Conversely, if the mask ratio is set too high, the [MASK] symbol may occur too frequently during pre-training, resulting in a mismatch between pre-training and fine-tuning, which could negatively impact performance. Hence, during the pre-training stage, a masked word ratio of $0.15$ is selected.
@@ 6.4.2. Masked Question Ratio
In our method, we design a question-level pre-training task for modeling expert interests. To explore the impacts of different masking question ratios on model performance, we varied the question masking ration in $[10\%,15\%,20\%,25\%,30\%]$ for experiments.
@@
Figure 8. Impacts of Different Historical Question Masking Ratios.
@@ 6.4.3. Learning Rate in Fine-tuning
During fine-tuning, we observe an interesting phenomenon: the learning rate significantly impacts expert finding performance. As depicted in Fig 7, we observe a consistent pattern across different datasets: the model performance initially improves and then gradually declines. A learning rate that is too large can potentially destroy the expert representations captured by the model during pre-training, whereas a learning rate that is too small may not capture sufficient information for identifying experts on a specific dataset. Additionally, the inflection point of the curve differs across different datasets. For instance, the optimal learning rate for the English dataset is 5e-8, while that of the Biology dataset is 5e-6. We think that this might be due to the size of the dataset. For larger datasets (e.g., English), a higher learning rate could lead to drastic changes in the pre-trained representations, resulting in degraded expert finding performance. Conversely, smaller datasets (e.g., Biology) may require a higher learning rate to achieve optimal performance. In summary, we recommend setting the learning rate carefully.
@@ 6.5. Case Study
In this part, we conduct a case study to better understand what has been learned by the PEPT during pre-training. Specifically, we randomly select several target questions from the English testing datasets and several pre-trained candidate experts for prediction. We visualize the final matching score $S_{c}$ in Fig 9, where the thickness of the arrow represents the matching score. A thicker arrow indicates a higher matching score between questions and candidate experts.
@@
These case studies demonstrate that our proposed paradigm can pre-train expert representations that include both interest and expertise, which can be transferred to the downstream task expert finding in CQA systems.
@@ 7. Conclusion
In this paper, our focus is on exploring the design of a personalized pre-training language model that is specifically customized for expert finding on CQA websites. Unlike the previous version ExpertPLM could only capture coarse-grained expert characteristics, our method is designed to precisely pre-train more personalized and fine-grained expert representations based on their histories. To achieve this goal, we propose a personalized pre-training framework that unifies expert historical answer questions, vote scores, and personalized information, then specially designs question-level masked language modeling and vote-oriented tasks. Based on these, the model is enabled to learn personalized expert features that can be transferred to the downstream task of expert finding, which could reduce the dependency on large amounts of labeled data. Our PEPT model demonstrates superiority expert finding performance on six real-world CQA datasets. Further ablation studies confirm the individual effectiveness of modeling interest, expertise, and personalization. Moreover, we conduct an evaluation to assess the performance of our model in scenarios with sparse training data. The results of this analysis highlight the model’s effectiveness in mitigating the issue of data sparsity.

### v2

@@ 5.4. Evaluation Metrics
For evaluating the model performance, we use the following three widely employed evaluation metrics in recommendation:
@@
Overall, the above metrics are used to assess the ranking accuracy of the expert recommended by each model. A higher value for each of these metrics indicates better performance of the model in accurately predicting the most relevant expert.
@@ 6. EXPERIMENTAL RESULTS
In this section, we represent the experiments and conduct a detailed analysis of PEPT. Specifically, (1) Performance Comparison compares the PEPT with baselines; (2) Ablation Study explores how different parts of the model affect model performance; (3) Pre-trained Expert Representation Analysis analyzes how the pre-trained expert representations benefit the expert finding. (4) Hyper-parameter Analysis analyzes the model performance with different model hyper-parameters. (5) Case Study understands PEPT’s advantages in finding experts more intuitively.
@@ 6.1. Performance Comparison
Table 4 shows the overall results of our proposed model in comparison to other baselines. The experimental results reflect following observations:
@@
    Early methods, such as CNTN, exhibit poor performance on almost all datasets. This attributes to the fact that these methods rely on the max or mean operator on experts’ histories to obtain expert features, which fails to capture the different importance of different historical questions in modeling the interest of an expert.
@@
    In contrast, recent attention-based methods, such as RMRN and PMEF, have achieved better results. The reason is that these methods leverage attention mechanisms to selectively weigh the importance of different historical questions in modeling expert representations. It is worth noting that the PMEF outperforms most of the baselines. This is because PMEF utilizes a multi-view attentive paradigm, enabling it to learn more comprehensive representations for experts and questions.
@@
    Our proposed method PEPT surpasses two very recent works ExpertBert (Liu et al., 2022b) and ExpertPLM (Peng and Liu, 2022). The reason is that the ExpertBert could not capture expert abilities to answer different questions and the ExpertPLM could not capture more fine-grained question-level interest and expertise. And both of them are incapable of modeling personalized expert representations, which might fail to accurately model different experts.
In summary, the results validate the motivation proposed in Section 1 and reflect that PEPT with all the characteristics in the Table 1 can achieve the best performance.
@@ 6.2. Ablation Study
Table 5. Characteristics of Different Variants. The variants possesses the feature marked with a check mark, while the feature marked with a cross is not possessed by the variant.
Characteristics Variants FT Bert Only Em w/o Expertise w/o Interest w/o Personalized PEPT
Expert-level Information $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$
Vote Score $\times$ $\times$ $\times$ $\surd$ $\surd$ $\surd$
ID Embedding $\times$ $\times$ $\surd$ $\surd$ $\times$ $\surd$
MLM $\times$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$
Vote-oriented Task $\times$ $\times$ $\times$ $\surd$ $\surd$ $\surd$
Question-level MLM $\times$ $\times$ $\surd$ $\times$ $\surd$ $\surd$
To further explore how different modules of the pre-training model impact the model performance, we design five variants for experimenting (CodeReview and Gis datasets as examples), including:
  1. (1)
    FT Bert, we directly fine-tune the original BERT (Devlin et al., 2018) on the CQA expert finding task, without further pre-training.
@@
    w/o Expertise, we incorporate expert interest modeling and personalized information during the pre-training. Excluding the vote score information and vote-oriented task from our model;
@@
A general comparison of the five variants is provided in Table 5, and the experimental results are illustrated in Fig 3. Based on these results, we have the following observations:
Figure 3. Ablation Study. We remove specific components in the PEPT during pre-training to evaluate their different effects on the overall performance. Detailed variants information can be found in Table 5.
@@
    Only Em obtains the better performance than FT Bert on the most metrics. This is because the Only Em uses expert-level word sequences (i.e., historical answered questions) to further pre-train the model, which could capture more CQA general knowledge for improving model performance.
@@
    w/o Expertise and w/o Interest show different degrees of decreased model performance. These results reflect the importance of capturing expert interest and capability for finding suitable experts. Interestingly, the performance of w/o Interest is worse than that of w/o Expertise. We argue that interest may be more important than ability in expert finding. In other words, experts in real life might first consider their interests rather than their abilities before answering questions;
@@
    Personalized information is crucial for modeling experts. Removing the personalized information during pre-training leads to a decrease in model performance, which can be seen in w/o Personalized;
@@
    Analysis of FT Bert. In fact, directly fine-tuning the BERT for expert finding can only find experts based on text similarity matching. From Table 5, we can find the characteristic difference between the model variants and the FT Bert. From Fig 3, we can find the performance of model variants is better than the FT Bert. Specifically: (1) When comparing the model without expertise information (w/o Expertise) and FT BERT, we can find w/o Expertise is superior to FT Bert. This is because the w/o Expertise can effectively capture the experts’ personalized fine-grained question-level interest information during pre-training; (2) Similarly, the model without interest information (w/o Interest) is also better that FT Bert. This phenomenon reveals that w/o Interest can successfully capture the experts’ personalized expertise information for the target question during pre-training; (3) Lastly, the comparison between the model without personalized information (w/o Personalized) and FT Bert shows that w/o Personalized is capable of simultaneously capturing both expert expertise and interest information during pre-training.
Overall, the ablation studies meet our motivation and demonstrate the effectiveness of each component in our proposed model.
@@ 6.3.1. Effect of Alleviating Data Sparsity
Figure 4. Impacts of Fine-tuning Data Ratio. The abscissa is the sparseness of the dataset, and the left ordinate is the MRR metric that measures model performance. GAP = PEPT - PMEF.
Fig 4 shows the results on CodeReview and Electronics datasets. The performance of all models substantially drops when a smaller amount of training data is used. However, our pre-trained model consistently outperforms the best non pre-trained model PMEF. Besides, we observe that with decreasing training data, the performance gap between PEPT and PMEF widens. This reflects that the advantage of our pre-trained representations becomes even more pronounced when the training data is limited. This is because our pre-training paradigm enables the model to personalized leverage expert behaviors and vote scores for modeling experts’ interests and expertise. Besides, it could reduce the model’s dependence on training data.
Furthermore, we can find that although ExpertPLM also has the ability to alleviate data sparsity issues, it is still inferior to PEPT. This is because ExpertPLM can only pre-train coarse-grained expert representations, and it still needs sufficient supervised data for fine-tuning to capture specific representations of different experts. Different from that, PEPT can learn personalized expert representations during pre-training and capture more fine-grained characteristics for expert finding. Overall, our experimental results validate that PEPT could mitigate the impact of the data sparsity problem in expert finding.
Figure 5. Zero-shot Performance.
Figure 6. Impacts of Different Word Masking Ratios.
6.3.2. Zero-shot Performance
In this part, we test the model performance with the zero-shot setting from Biology to Es and Electronics (we fine-tune the model based on the Biology dataset, and test the performance on the Es and Electronics). The experimental results are shown in Fig 5.
It is surprising that the model can still achieve good performance (especially in P@3) by simply transferring from Biology to Es/Electronics without any fine-tuning. The reason could be that the model has already captured a large amount of expert and question related knowledge during pre-training, and merely fine-tuning on smaller datasets (e.g., biology) could stimulate model understanding of the expert finding task, leading to zero-shot performance on other large datasets.
@@ 6.4. Further Analysis
In this section, we conduct experiments to evaluate the effect of key hyperparameters (including Masked Word Ratio, Masked Question Ratio, Learning Rate in Fine-tuning) in our approach.
@@ 6.4.1. Masked Word Ratio
The masked word ratio is a hyperparameter in our proposed PEPT model during the expert representation pre-training. To investigate the sensitivity of that in the pre-training, we vary the ratio from $0.05$ to $0.25$ in increments of $0.05$.
The results are presented in Fig 6. All metric results initially improve as the mask ratio increases, reaching the optimal model performance at a certain point, and then gradually declining. If the word mask ratio is set too small, the pre-training process may not capture enough CQA language knowledge and result in reduced performance in the downstream task. A high mask ratio during pre-training may lead to excessive [MASK] symbol occurrence, causing a mismatch and negatively impacting performance during fine-tuning. Hence, during the pre-training stage, a masked word ratio of $0.15$ is selected.
@@ 6.4.2. Masked Question Ratio
In our method, we design a question-level pre-training task for modeling expert interests. To explore the impacts of different masking question ratios on model performance, we vary the question masking ration in $[10\%,15\%,20\%,25\%,30\%]$ for experiments.
Figure 7. Impacts of Different Historical Question Masking Ratios.
@@ 6.4.3. Learning Rate in Fine-tuning
During fine-tuning, we observe an interesting phenomenon: the learning rate significantly impacts expert finding performance. As depicted in Fig 8, we observe a consistent pattern across different datasets: the model performance initially improves and then gradually declines. A large learning rate can harm expert representations from pre-training. Additionally, the inflection point of the curve differs across different datasets. For instance, the optimal learning rate for the English dataset is 5e-8, while that of the Biology dataset is 5e-6. We think that this might be due to the size of the dataset. For larger datasets (e.g., English), a higher learning rate could lead to drastic changes in the pre-trained representations, resulting in degraded expert finding performance. Conversely, smaller datasets (e.g., Biology) may require a higher learning rate to achieve optimal performance. In summary, we recommend setting the learning rate carefully.
@@ 6.5. Case Study
In this part, we conduct a case study to better understand what has been learned by the PEPT during pre-training. Specifically, we randomly select several target questions from the English testing datasets and several pre-trained candidate experts for prediction. We visualize the final matching score $E_{c}$ in Fig 9, where the thickness of the arrow represents the matching score. A thicker arrow indicates a higher matching score between questions and candidate experts.
@@
These case studies demonstrate that our paradigm pre-trains expert representations with both interest and expertise, benefiting expert finding in CQA.
7. Limitation
In fact, the vote scores received for each question can roughly measure the answerer’s expertise in the question. However, because the popularity of questions varies, a low vote score may not necessarily mean that the answer is of poor quality. For alleviating this, in the future, we will explore the following aspects:
  1. (1)
    Categorizing different questions and adjusting the weight of vote scores based on the category popularity. For example, questions in less popular areas usually have fewer votes, so the weight of such question vote scores can be amplified.
  2. (2)
    Mitigating bias caused by question popularity from a causal perspective. First, we can construct a causal graph to describe the relationship between questions, answers, voting scores, popularity and categories. The relationship might include: category affects question, question affects answer, answer affects voting score, popularity of a question affects the question, and user expertise affects answer, etc.
    In this causal model, the popularity of the question has a indirect impact on the voting score, which may lead to a deviation in the voting scores of questions of different categories. Then we employ inverse probability weighting to adjust the voting score of each question, so that the voting score can more reflect the user’s true professional level.
  3. (3)
    Introducing Large Language Models (LLMs, e.g, ChatGPT) to evaluate the professionalism of the expert’s historical answers because the LLMs has large general knowledge. Specifically, we fine-tune the open source LLMs (e.g., Llama-3) or directly use the closed source LLMs (e.g., ChatGPT) to evaluate the expert expertise. Given one expert, we input historical question and corresponding answer into the LLMs, and employ the prompt (e.g., You are a CQA expert. I have given you a question and a corresponding answer. Please help me judge the professionalism of the answer (score: 1-5).) to make the model output a professional score. Then we introduce this score into the expert finding model as auxiliary information to help model expert ability.
@@ 8. Conclusion
In this paper, our focus is on exploring the design of a personalized pre-training language model that is specifically customized for expert finding on CQA websites. Unlike the previous version ExpertPLM could only capture coarse-grained expert characteristics, we propose a more fine-grained pre-training framework that integrates expert histories, vote scores, and personalized information. This framework includes question-level masked language modeling and vote-oriented tasks. Based on these, the model is enabled to learn personalized expert features that can be transferred to the downstream task of expert finding, which could reduce the dependency on large amounts of labeled data. Our PEPT model demonstrates superiority expert finding performance on six real-world CQA datasets. Further ablation studies confirm the individual effectiveness of modeling interest, expertise, and personalization. Moreover, we conduct an evaluation to assess the performance of our model in zero-shot scenarios, which could validate that the expert personalized information captured by the model during pre-training.
Acknowledgements.
This work was supported by: (1) Hainan Province Science and Technology Special Fund, Grant No: ZDYF2024SHFZ051; (2) The Research Program of Yazhou Bay Innovation Research Institute, Hainan Tropical Ocean University, Grant No: 2023CXYZD001.